# Tony Lindeberg

### v2

@@ 2.3 Orientation selectivity curves for the idealized receptive field models
Based on these results we propose that, beyond an expansion over rotations, as is performed in current models of the pinwheel structure of visual receptive fields (Bonhoeffer and Grinvald 1991; Blasdel 1992; Swindale 1996; Petitot 2003; Baspinar et al. 2018; Liu and Robinson 2022), also an explicit expansion over the eccentricity $\epsilon$ of the receptive fields (the inverse of the parameter $\kappa$) should be included, when modelling the pinwheel structure in the visual cortex.
@@
4 Explicit testable hypotheses for biological experiments
Based on the above theoretical analysis, with its associated theoretical predictions, we propose that it would be highly interesting to perform experimental characterization and analysis based on joint estimation of (i) orientational selectivity, (ii) receptive field eccentricity, (iii) orientational homogeneity and (iv) location of the neuron in the visual cortex in relation to the pinwheel structure, in the primary visual cortex of animals with clear pinwheel structures, to determine if there is a variability in the eccentricity or elongation of the receptive fields, and specifically if the degree of elongation increases with the distance from the centres of the pinwheels towards periphery, as arising as one possible interpretation of combining the theoretical results about orientation selectivity of affine Gaussian receptive fields in this article with the biological results by Nauhaus et al. (2008).
If additionally, reconstructions of the receptive field shapes could be performed for the receptive fields probed during such a systematic investigation of the difference in response characteristics with the distance from the pinwheel centers, and if the receptive fields could additionally be reasonably well modelled according to the generalized Gaussian model for receptive fields studied and used in this paper, it would be interesting to investigate if the shapes of the affine Gaussian components of these receptive fields would span a larger part of the affine group, than the span over mere image orientations, as already established in the orientation maps of the visual cortex, as characterized by Bonhoeffer and Grinvald 1991, Blasdel 1992 and others, see Figure 7 for an illustration.
@@ 4 Explicit testable hypotheses for biological experiments
Notably the variability of the spatial covariance matrices in the affine Gaussian derivative model comprises a variability over two spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$, while the theoretical analysis of the orientation selectivity properties studied in this article has mainly concerned their ratio $\kappa=\sigma_{2}/\sigma_{1}$. Hence, the illustration in Figure 7 should not be taken as a literal prediction, even if reduced by a double-angle representation. In Figure 7, the larger scale parameter $\sigma_{1}$ is held constant, for convenience of graphical illustration, as obtained by mapping the uniformly sized receptive fields from a uniform distribution on the hemisphere. More generally, one could also conceive other distributions as possible, such as, for example, instead keeping the smaller eigenvalue $\sigma_{2}$ constant from the center towards the periphery.
@@
It should finally be stressed that, we have in this treatment not considered the binocular aspects of the pinwheel structure. In Hypothesis 5, the variability of the pinwheel structure over contributions from the left and the right eyes should therefore not be counted as a property to contribute to the terminology “more structured”.
5 Quantitative measurements for detailed characterization
@@
6 Relations to previous work
Beyond the work by Nauhaus et al. (2008), which the treatment in this paper largely builds upon, there is a large body of work on characterizing the orientation selectivity of neurons, by Rose and Blakemore (1974), Schiller et al. (1976), Ringach et al. (2002), Nauhaus et al. (2008), Scholl et al. (2013), Sadeh and Rotter (2014), Goris et al. (2015) and Sasaki et al. (2015), as well as concerning biological mechanisms for achieving orientation selectivity by Somers et al. 1995, Sompolinsky and Shapley 1997, Carandini and Ringach 1997, Lampl et al. 2001, Ferster and Miller 2000, Shapley et al. 2003, Seriès et al. 2004 and Priebe (2016). The focus of this paper work is, however, not on the neural mechanisms that lead to orientation selectivity, but concerning purely functional properties at the macroscopic level.
Mathematical models of biological receptive fields have beyond in terms of Gaussian derivatives (Koenderink and van Doorn 1984; 1987; 1992; Young and his co-workers 1987; 2001; 2001, Lindeberg 2013a; 2021) also been formulated in terms of Gabor filters (Marcelja 1980; Jones and Palmer 1987a; 1987b; Porat and Zeevi 1988). Gaussian derivatives have, in turn, been used as primitives in theoretical models of early visual processing by Lowe (2000), May and Georgeson (2007), Hesse and Georgeson (2005), Georgeson et al. (2007), Wallis and Georgeson (2009), Hansen and Neumann (2008), Wang and Spratling (2016) and Pei et al. (2016).
Hubel and Wiesel (1959; 1962; 2005) pioneered the study of simple and complex cells. The properties of simple cells have been further characterized by DeAngelis et al. (1995; 2004), Ringach (2002; 2004), Conway and Livingstone (2006), Johnson et al. (2008) and De and Horwitz (2021), and the properties of complex cells investigated by Movshon et al. (1978), Emerson et al. (1987), Touryan et al. (2002; 2005), Rust et al. (2005) and Goris et al. (2015), as well as modelled computationally by Adelson and Bergen (1985), Heeger (1992), Serre and Riesenhuber (2004), Einhäuser et al. (2002), Kording et al. (2004), Merolla and Boahen (2004), Berkes and Wiscott (2005), Carandini (2006) and Hansard and Horaud (2011). In this work, we have followed a specific way of modelling simple and complex cells in terms of affine Gaussian derivatives, according to the generalized affine Gaussian derivative model for visual receptive fields.
7 Summary and discussion
Based on a combination of theoretical results, concerning a direct connection between the degree of orientation selectivity and the degree of elongation of receptive fields, in combination with the results from biological experiments by Nauhaus et al. (2008), which show a substantial variability in the degree of orientation selectivity of the neurons in the primary visual cortex, we have addressed the question of whether the receptive fields in the primary visual cortex can be regarding as spanning a variability in the degree of elongation.
An in-depth theoretical analysis in a companion paper (Lindeberg 2024), as summarized in Section 2, shows how the orientation selectivity depends on a scale parameter ratio $\kappa$, between the scale parameters in the image orientations perpendicular to vs. parallel with the preferred orientation of the receptive fields, for the generalized Gaussian derivative model for visual receptive fields. The explicit expressions for the resulting orientation tuning curves, derived from closed-form theoretical analysis, show that, for both (i) models of simple cells in terms of first-order affine Gaussian derivatives, (ii) models of simple cells in terms of second-order affine Gaussian derivatives, and (iii) models of complex cells in terms of quasi-quadrature measures of simple cells in terms of first- and second-order Gaussian derivatives, the orientation selectivity becomes more narrow with increasing values of the scale parameter ratio $\kappa$. This result holds for both purely spatial models of the receptive fields as well as for joint spatio-temporal receptive field models.
By comparisons with previously established results by Nauhaus et al. (2008), we have in Section 3 demonstrated that the predictions from our theory are consistent with general behaviour of visual neurons in the primary visual cortex, in that the orientation tuning is broader near the pinwheels, while the orientation selectivity becomes sharper further away.
@@ 7 Summary and discussion
By combining these two results, if we, would take the liberty of assuming that it would be unlikely for the population of receptive fields to exhibit a strong variability in the orientational selectivity of the receptive fields, without also similarly exhibiting a coupled variability in the eccentricity or the the degree of elongation of the receptive fields, then we would have possible indirect support for a previously formulated biological hypothesis in (Lindeberg 2023), stating that the family of receptive field shapes should span the degrees of freedom in the natural geometric image transformations. That way of assumption-based reasoning would then specifically imply indirect support for the hypothesis that the receptive field shapes should span a sufficiently wide range of ratios between the scale parameters in the directions perpendicular to vs. parallel with the preferred orientation of the receptive field, to support affine covariance of the family of visual receptive fields.
@@
Without explicitly relying on expressing such an explicit assumption, regarding whether the visual receptive fields in the primary visual cortex could be well modelled by affine Gaussian derivative based receptive fields, we can, however, firmly state that the biological measurements performed by Nauhaus et al. (2008) are, in combination with the theoretical results summarized in Section 2, consistent with the hypothesis that the receptive fields should span a variability in the eccentricity of the receptive fields. In this respect, the measurements that demonstrate a strong variability in orientation selectivity would specifically be consistent with the theoretically based hypothesis formulated in (Lindeberg 2023), that the receptive fields in the primary visual cortex should span the variability of receptive field shapes under spatial affine transformations.
@@
If we would apply a similar type of assumption-based logical reasoning to the pinwheel structure in the primary visual cortex, then such a reasoning, based on the results by Nauhaus et al. (2008), that the orientation selectivity appears to vary strongly from the centers of the pinwheels towards the periphery, would imply that the pinwheel structure in the visual cortex would, beyond an explicit expansion over image orientations, also comprise an explicit expansion over the eccentricity or the degree of elongation of the receptive fields. Based on these predictions, we have specifically proposed that explicit dependencies on a variability in the eccentricity of the receptive fields should be included, when modelling the pinwheel structures in the primary visual cortex.
@@
Strictly, and formally, the results from such logical inference could, however, only be regarded as theoretical predictions, to generate explicit hypothesis concerning the distribution of receptive field characteristics in these respects. To raise the question of determining if these theoretical predictions would hold in reality, we have in Section 4 formulated a set of testable explicit biological hypotheses, that could be either verified or rejected in neurophysiological experiments, concerning possible variabilities in the eccentricity of the receptive fields in the primary visual cortex of higher mammals, as well as hypotheses about possible connections between such variabilities in the eccentricity or the elongation and other receptive field characteristics, in particular in relation to the pinwheel structure in the primary visual cortex of higher mammals. We have also in Section 5 formulated a set of quantitative measurements that could be made, to characterize how a possible variability in the eccentricity of the receptive field could be related to other receptive field characteristics, including the pinwheel structure in the visual cortex.

### v7

@@ 3 Results
Based on these results we propose that, beyond an expansion over rotations, as is performed in current models of the pinwheel structure of visual receptive fields (Bonhoeffer and Grinvald 1991, Blasdel 1992, Swindale 1996, Petitot 2003, Koch et al. 2016, Kremkow et al. 2016, Baspinar et al. 2018, Najafian et al. 2022, Liu and Robinson 2022), also an explicit expansion over the eccentricity $\epsilon$ of the receptive fields (the inverse of the parameter $\kappa$) should be included, when modelling the pinwheel structure in the visual cortex.
@@
3.2 Explicit testable hypotheses for biological experiments
Based on the above theoretical analysis, with its associated theoretical predictions, we propose that it would be highly interesting to perform experimental characterization and analysis based on joint estimation of
  • orientational selectivity,
  • receptive field eccentricity,
  • orientational homogeneity and
  • location of the neuron in the visual cortex in relation to the pinwheel structure,
in the primary visual cortex of animals with clear pinwheel structures, to determine if there is a variability in the eccentricity or elongation of the receptive fields, and specifically if the degree of elongation increases with the distance from the centres of the pinwheels towards periphery, as arising as one possible interpretation of combining the theoretical results about orientation selectivity of affine Gaussian receptive fields in this article with the biological results by Nauhaus et al. (2008).
If additionally, reconstructions of the receptive field shapes could be performed for the receptive fields probed during such a systematic investigation of the difference in response characteristics with the distance from the pinwheel centers, and if the receptive fields could additionally be reasonably well modelled according to the generalized Gaussian model for receptive fields studied and used in this paper, it would be interesting to investigate if the shapes of the affine Gaussian components of these receptive fields would span a larger part of the affine group, than the span over mere image orientations, as already established in the orientation maps of the visual cortex, as characterized by Bonhoeffer and Grinvald (1991), Blasdel (1992) and others, see Figure 13 for an illustration.
@@ 3.2 Explicit testable hypotheses for biological experiments
Notably the variability of the spatial covariance matrices in the affine Gaussian derivative model comprises a variability over two spatial scale parameters $\sigma_{1}$ and $\sigma_{2}$, while the theoretical analysis of the orientation selectivity properties studied in this article has mainly concerned their ratio $\kappa=\sigma_{2}/\sigma_{1}$. Hence, the illustration in Figure 13 should not be taken as a literal prediction, even if reduced by a double-angle representation. In Figure 13, the larger scale parameter $\sigma_{1}$ is held constant, for convenience of graphical illustration, as obtained by mapping the uniformly sized receptive fields from a uniform distribution on the hemisphere. More generally, one could also conceive other distributions as possible, such as instead keeping the smaller eigenvalue $\sigma_{2}$ constant from the center towards the periphery.
@@
It should finally be stressed that, in this treatment, we have not considered the binocular aspects of the pinwheel structure. In Hypothesis 5, the variability of the pinwheel structure over contributions from the left and the right eyes should therefore not be counted as a property to contribute to the terminology “more structured”.
3.3 Quantitative measurements for detailed characterization
@@
4 Discussion
By comparing the theoretical analysis of the orientation selectivity properties of the affine Gaussian derivative model (Section 2.2) with the experimental results from Nauhaus et al. (2008) on broadly vs. sharply tuned visual neurons (Figure 6) and from Goris et al. (2015) on rather uniform distributions of the resultant values from orientation selectivity curves (Figure 7), we find potential support for one of the dimensions of variability in a biological hypothesis formulated in Lindeberg (2023), stating that the family of receptive field shapes should span the degrees of freedom in the natural geometric image transformations. This potential support rests on the assumption, that it should be unlikely for the population of receptive fields to show strong variability in orientation selectivity, without also showing similar variability in eccentricity or elongation.
@@ 4 Discussion
Such an assumption-based reasoning would then specifically imply indirect support for the hypothesis that the receptive field shapes should span a sufficiently wide range of ratios between the scale parameters in the directions perpendicular to versus parallel with the preferred orientation of the receptive field, to support affine covariance of the family of visual receptive fields.
@@
Without explicitly relying on expressing such an explicit assumption, regarding whether the visual receptive fields in the primary visual cortex could be well modelled by affine Gaussian derivative based receptive fields, we can, however, firmly state that the biological measurements performed by Nauhaus et al. (2008) and by Goris et al. (2015) are, in combination with the theoretical results summarized in Section 2.2, consistent with the hypothesis that the receptive fields should span a variability in the eccentricity of the receptive fields. In this respect, the measurements that demonstrate a strong variability in orientation selectivity would specifically be consistent with the theoretically based biological hypothesis formulated in (Lindeberg 2023), that the receptive fields in the primary visual cortex should span the variability of receptive field shapes under spatial affine transformations.
Furthermore, from the results presented in Section 3.1.3, that the orientation selectivity histograms accumulated by Goris et al. (2015 appear to be better explained by a population of receptive fields up to order 4 than by receptive fields up to order 2, one may speculate if it would be appropriate to put additional focus of neurophysiological experiments on exploring the presence of visual receptive fields of higher order, compared to the receptive fields of lower order, that are more commonly reported in the neurophysiological literature.
@@
If we would apply a similar type of assumption-based logical reasoning to the pinwheel structure in the primary visual cortex, then such a reasoning, based on the results by Nauhaus et al. (2008), that the orientation selectivity appears to vary strongly from the centers of the pinwheels towards the periphery, would imply that the pinwheel structure in the visual cortex would, beyond an explicit expansion over image orientations, also comprise an explicit expansion over the eccentricity or the degree of elongation of the receptive fields. Based on these predictions, we propose to consider explicit dependencies on a variability in the eccentricity of the receptive fields, when modelling the pinwheel structures in the primary visual cortex.
@@
Strictly, and formally, the results from such logical inference could, however, only be regarded as theoretical predictions, to generate explicit hypothesis concerning the distribution of receptive field characteristics in these respects. To raise the question of determining if these theoretical predictions would firmly hold in reality, we propose that the testable explicit biological hypotheses formulated in Section 3.2 could be used to, in neurophysiological experiments, either verify or reject the overall hypothesis, concerning possible variabilities in the eccentricity of the receptive fields in the primary visual cortex of higher mammals, as well as hypotheses about possible connections between such variabilities in the eccentricity or the elongation and other receptive field characteristics, in particular in relation to the pinwheel structure in the primary visual cortex of higher mammals. Furthermore, if those hypothesis would hold, then the proposed quantitative measurements formulated in Section 3.3 could be used, to characterise how a possible variability in the eccentricity of the receptive field could be related to other receptive field characteristics, including the pinwheel structure in the visual cortex.